mario wants to rent a ride a bike at a state park. there are 2 parks in his area. one has an entrance fee of $8 and charges $2 per hour for bike rentals. the other park has an entrance fee of $2 and charges $5 an hour for bike rentals. After riding how many hours will the cost be the same? what would the cost be?
Answer by checkley77(12844)

8+2X=2+5X
2X-5X=2-8
-3X=-6
X=-6/-3
X=2 HOURS RIDING WILL MAKE THE TWO RENTALS THE SAME.
PROOF:
8+2*2=2+5*2
8+4=2+10
12=12